OpenStudy (anonymous):

List the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1.

OpenStudy (anonymous):

\(t_n= -3t_{n-1}\), right?

OpenStudy (campbell_st):

so for the 2nd term \[t_{n -1} = t_{1}\] so the 2nd term is -3 x 4 the 3rd term is -3 x the 2nd term 4th term = -3 x 3rd term

OpenStudy (anonymous):

-12 36 -108 324

OpenStudy (campbell_st):

well you were given the 1st tem... and asked to find the 1st 4... so you only need the next 3 4, -12, 36, -108

OpenStudy (anonymous):

Find the sixth term of a geometric sequence with t5 = 24 and t8 = 3. how would i do this

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well a term in a geometric sequence is \[t_{n} = ar^{n - 1}\] so you need to find the common ratio r.... a is the 1st term so using the known information \[t_{5} = a \times r^4 = 24~~~~t_{8} = a \times r^7 = (a \times r^4) \times r^3\] so substituting \[3 = 24 \times r^3\] find r.... then the 6th term is the 5th term times the value of r